OpenStudy (anonymous):

Help!!!! How do I integrate this: integral (2x)/(sqrt1-x^2) from 0 to 1/2

OpenStudy (anonymous):

\[\int\limits_{0}^{1/2}2x/\sqrt{1-x^2} dx\]

OpenStudy (anonymous):

try \(u=1-x^2\)

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{ 1 }{ 2 }}\frac{ 2x }{ \sqrt{1-x^2} }dx\] put \[x^2=t,2 x dx=dt\] when x=0,t=o \[when~ x=\frac{ 1 }{ 2 },t=\frac{ 1 }{ 4 }\] can you proceed further?

OpenStudy (anonymous):

no I have no idea how to do this

OpenStudy (anonymous):

you write the following: "let \(u=1-x^2\) and so \(du=-2xdx\)

OpenStudy (anonymous):

then take your integral and replace \(1-x^2\) by \(u\) and replace the \(2xdx\) by \(-du\) pull the minus sign out front

OpenStudy (anonymous):

that gives you \[-\int \frac{du}{\sqrt{u}}\]

OpenStudy (anonymous):

now find the anti derivative by writing this in exponential form and using the power rule backwards